# Benchmarking Deep Learning-Based Methods for Irradiance Nowcasting with Sky Images

### v1

@@ 3.6 Sky Image and Irradiance Alignment
[…]
Figure 5: A 5-minute GHI window in $\mathcal{D}^{\text{folsom}}$ showing the corresponding date modified aligned images in the top row and the file name aligned images in the bottom row. The middle row shows the linearly interpolated GHI data, plotted in UTC-8 timezone for better visualization, while the original file names (titles above each image) are in UTC+0 timezone.
In the 5-minute window being shown in Figure 5, the file name is approximately 11:30 minutes ahead of the date modified, but, from our analysis, we found that this difference is shifting throughout the dataset, which can be seen in Figure 6, that shows the daily average difference between the file name timestamp and the date modified timestamp. A positive value means that the file name is ahead of the date modified and vice-versa. As time passes, this difference seem to be shifting towards 0 seconds, which would be ideal, however 3 major jumps occur in Jun-2014, Feb-2015 and Sep-2016, which significantly aggravate the problem, especially towards the end of 2016, where a 12-minute difference can be observed between the two. This means that, assuming the date modified is the correct timestamps of the images (Figure 5 indicates that this is a fair assumption), then aligning the irradiance data with the images based on their file names would always result in incorrect labels for the model to learn, as well as for model evaluation. Therefore, we use the date modified alignment in $\mathcal{D}^{\text{folsom}}$, except during the experiment in subsubsection 5.2.1, where we provide a more quantitative analysis on this issue.
@@
As a precautionary measure, we also compared the metadata of $x^{\text{sirta}}$ and $x^{\text{nrel}}$ to their file names, but no inconsistency was found. Therefore, the sky image and irradiance alignment in those two datasets was done based on the image file names, as that is the most natural choice.
Another thing to keep in mind when aligning the images with the irradiance data is the irregularity in the image timestamps, which is especially true assuming the date modified as the correct timestamp (see Figure 7). Since the irradiance samples are all synchronized at :00 seconds, aligning an image that was captured at :30 seconds with the nearest irradiance sample could result in a large error, especially in fast moving cloud conditions. To mitigate this problem, we linearly interpolate the irradiance samples before the alignment. While this mostly affects $\mathcal{D}^{\text{folsom}}$, the linear interpolation is applied to the other two datasets as well.
Figure 7: Image timestamp second counts in $\mathcal{D}^{\text{folsom}}$. The upper axis shows that the file name timestamps are mostly synchronized at either :00 or :59 seconds (i.e. the file ends in “00.jpg” or “59.jpg”). On the other hand, the lower axis shows that the date modified timestamps are uniformly spaced throughout the 60 possible values.
@@ 3.7 Sun Mask Generation
In subsubsection 5.2.6 we conduct an ablation experiment where we use a Sun mask as an additional input to the nowcasting model, which has shown to significantly benefit the performance in [22]. First, we compute the distance of the Sun center from the center of the image plane, denoted by R, which is a function of the focal length of the camera f and the angle from the optical axis $\Phi$. This mapping function will vary, since different types of lenses will have different projection types [52]. For the stereographic projection, the mapping function is $\text{R}=2f\tan{\frac{\Phi}{2}}$, and for the equidistant projection the mapping function is $\text{R}=f\Phi$, where $\Phi$ is replaced by the zenith angle $\theta_{z}$. Following this, we compute the Sun center coordinates on the image plane according to Equation 7 and Equation 8.
@@ 3.8 Train, Test and Validation Splits
The remaining years of each data set are used for the train set. In order to optimize the hyperparameters of the model, we split the full train set into a smaller train set and a validation set using $K$-Fold cross validation (CV). More specifically, we use stratified group $K$-Fold (with $K=5$), where each day is assigned to a different group (to avoid leaking samples from the same day in the training/validation sets) and each bin in the original train set (see Figure 3) is assigned to a different class to maintain the original irradiance distribution in all 5 folds. After optimizing the hyperparameters of each model, we use the full train set to train a final model, which will be evaluated on the test set. Table 2 shows the number of samples in each set, after the application of all the pre-processing steps.
Table 2: Number of samples for each set after the application of all the pre-processing steps.
Test
Dataset Train All Clear Cloudy
Folsom 426,515 220,755 117,423 103,332
SIRTA 297,774 103,926 18,447 85,479
NREL 59,444 20,101 7,300 12,801
@@ 4 Methods
In this section we go into detail about our training setup, the model architectures we trained and our evaluation metrics. As the main of objective of this study is to benchmark different model architectures for the on-site irradiance nowcasting task, all of our models are trained and evaluated separately for each dataset, however we also evaluate the aggregated performance across all three datasets.
4.1 Training Setup
Given an input image $x_{i}$, each model produces an estimate $\hat{I}_{i}$ of the concurrent GHI measurement $I_{i}$. As described in subsection 3.6, the sky image and irradiance measurement pairs $\{x_{i},I_{i}\}$ are created using the image file name timestamps in $\mathcal{D}^{\text{sirta}}$ and $\mathcal{D}^{\text{nrel}}$ and the image date modified timestamps in $\mathcal{D}^{\text{folsom}}$. The models are trained to minimize the Mean Squared Error (MSE) loss function given by
@@ 4.1 Training Setup
Our models are trained with a learning rate that exponentially decays from its initial value $\text{lr}_{0}$ to $\frac{\text{lr}_{0}}{10}$ during the first 75% of training according to Equation 12
@@
[…]
4.2 Model Architectures and Hyperparameter Optimization
@@
We search 30 different hyperparameter combinations per model per dataset, totaling 900 different searches, which are made through Bayesian optimization [57] using the train/validation splits defined in subsection 3.8 and with the goal of minimizing the average validation loss across all 5 folds. For hyperparameter optimization, we reduce the number of training epochs to 8 and only use 10% of the training data in each dataset. The hyperparameter search space is shown in Table 3.
@@
Initial Learning Rate ($\text{lr}_{0}$) $[1,2,3,4,5,6,7,8,9,10]\times 10^{-4}$
@@
4.3 Evaluation Metrics
A large variety of metrics have been proposed by the solar forecasting community in order to assess the quality of a model’s predictions in the most insightful way possible [60, 61, 62]. In this study, we only evaluate the root mean squared error (RMSE) and the mean absolute error (MAE), which are well established metrics for evaluating the overall fit of the predictions:
@@ 4.3 Evaluation Metrics
[…]
@@ 5 Experiments and Results
5.1 Model Architectures
Table 4 shows the RMSE and MAE for the test set for all ten models and all three datasets, as well as the aggregated performance across all datasets. From these results, it is clear that the ResNet50 architecture outperforms the others by a healthy margin, especially in comparison to the Sunset and UNet architectures. The only 2 cases where the ResNet50 was not the best performing architecture was in cloudy sky conditions in $\mathcal{D}^{\text{folsom}}$ (where the EfficientNetV2-S performed better) and in clear sky conditions in $\mathcal{D}^{\text{nrel}}$ (where the VGG16 performed better).
@@ 5.1 Model Architectures
NREL EfficientNet 63.40 ± 0.53 78.95 ± 0.60 11.63 ± 1.05 32.22 ± 0.50 45.71 ± 0.29 8.58 ± 0.97
MobileNetV2 64.81 ± 0.34 80.58 ± 0.45 13.43 ± 1.32 33.73 ± 0.53 47.35 ± 0.35 9.85 ± 1.19
RegNet 65.06 ± 0.59 81.03 ± 0.72 12.00 ± 0.67 32.94 ± 0.38 46.63 ± 0.21 8.93 ± 0.74
ResNet18 102.68 ± 45.09 104.77 ± 42.09 79.30 ± 77.36 37.91 ± 5.03 49.94 ± 5.55 16.81 ± 8.69
ResNet34 65.28 ± 0.99 81.08 ± 1.23 13.90 ± 3.37 33.09 ± 0.35 46.85 ± 0.57 8.96 ± 0.35
ResNet50 62.10 ± 0.37 77.31 ± 0.50 11.63 ± 0.82 31.47 ± 0.32 44.54 ± 0.34 8.54 ± 0.72
SolarNet 67.22 ± 0.80 83.67 ± 1.01 12.88 ± 0.45 35.05 ± 0.53 49.37 ± 0.77 9.93 ± 0.41
Sunset 92.07 ± 2.15 113.58 ± 2.68 26.81 ± 1.42 55.14 ± 1.13 76.44 ± 1.59 17.79 ± 1.03
UNet 71.34 ± 2.76 88.50 ± 3.43 16.58 ± 1.88 39.38 ± 1.91 55.15 ± 2.74 11.73 ± 1.13
VGG16 64.68 ± 0.44 80.64 ± 0.55 10.81 ± 0.15 32.69 ± 0.30 46.86 ± 0.44 7.86 ± 0.08
Aggr. EfficientNet 40.20 ± 0.06 51.33 ± 0.09 13.47 ± 0.44 22.39 ± 0.16 31.14 ± 0.14 10.07 ± 0.29
MobileNetV2 45.27 ± 1.34 56.87 ± 1.23 19.38 ± 2.74 26.96 ± 1.28 35.69 ± 1.01 14.66 ± 1.89
RegNet 42.11 ± 0.82 53.74 ± 0.85 14.25 ± 1.37 23.74 ± 0.71 33.15 ± 0.68 10.50 ± 0.79
ResNet18 48.35 ± 6.32 58.32 ± 6.39 26.39 ± 13.53 25.30 ± 0.72 34.41 ± 1.02 12.47 ± 0.53
ResNet34 46.96 ± 7.61 53.40 ± 0.33 31.19 ± 21.47 24.60 ± 1.18 32.35 ± 0.30 13.69 ± 2.88
ResNet50 39.82 ± 0.21 50.99 ± 0.24 12.57 ± 0.36 21.96 ± 0.15 30.80 ± 0.18 9.51 ± 0.22
SolarNet 44.93 ± 1.29 56.74 ± 1.14 17.95 ± 2.57 25.86 ± 1.03 34.82 ± 0.77 13.24 ± 1.42
Sunset 51.38 ± 1.20 64.12 ± 1.28 23.67 ± 2.99 30.77 ± 1.20 40.23 ± 0.83 17.44 ± 2.27
UNet 50.94 ± 1.39 61.10 ± 1.01 31.23 ± 4.38 33.04 ± 1.78 39.72 ± 0.85 23.62 ± 3.55
VGG16 45.66 ± 0.97 57.64 ± 0.99 18.46 ± 1.41 26.72 ± 0.74 36.44 ± 0.73 13.03 ± 0.76

### v2

@@ 3.6 Sky Image and Irradiance Alignment
Our sky image and irradiance alignment for each dataset was very different, especially for $\mathcal{D}^{\text{folsom}}$. Initially, we aligned the sky images with the irradiance data based on the name of the image files, i.e. an image sample $x^{\text{folsom}}_{t}$, with file name timestamp $t$, was labeled with the irradiance sample $I^{\text{folsom}}_{t}$. While this seems like the most natural choice, we noticed, during our initial experiments, that the model’s largest errors were samples were the prediction actually seemed quite reasonable, since it predicted larger GHI values when the Sun was clearly visible and lower values when it was not. This raised the suspicion that the labels might be misaligned. To the best of our knowledge, this unexpected behavior in $\mathcal{D}^{\text{folsom}}$ has not yet been discussed in the literature and, upon further investigation, we found that the samples taken during time windows with fast moving cloud conditions (e.g. Figure 6) corresponded to the highest errors, leading us to discover a discrepancy between the date modified and the image file name timestamps. In Figure 6 we show two distinct 5 minute GHI/DNI windows in the middle row, with the image date modified alignment in the top row and the image file name alignment in the bottom row. Both cases show that the GHI began at a medium-to-low value, which indicates some cloud coverage, before rapidly increasing, which indicates loss of cloud coverage and a more visible Sun. This behavior seems much more consistent with the top row images, strongly suggesting that it is actually the date modified timestamps that correctly correspond to the GHI time series. However, due to the strong presence of clouds, one could argue that this increase in GHI could be due to an extreme overirradiance event. However, since the DNI component is also being plotted, we can rule out this possibility and confirm that the Sun should indeed transition from covered to visible during both time windows.
Figure 6: Two distinct 5-minute GHI and DNI windows in $\mathcal{D}^{\text{folsom}}$ which show the corresponding Date Modified (DM) aligned images in the top row and the File Name (FN) aligned images in the bottom row. The middle row shows the linearly interpolated GHI data, plotted in UTC-8 timezone. The titles above each image indicate their corresponding FN and DM timestamps (converted to UTC-8 timezone), as well as their JPEG file names. The * marker indicates the timestamp used to align the image to the irradiance plot.
In the 5-minute windows being shown in Figure 6, the file name is approximately 11 minutes 30 seconds ahead of the date modified, but, from our analysis, we found that this difference changes throughout the dataset, as depicted in Figure 7, which clearly indicates a time drift between both timestamps [^1] We hypothesize that the file name timestamps are based on the camera’s internal clock, while the date modified is based on the database’s clock. Assuming the irradiance timestamps are synchronized with the latter, then the file name timestamp would indeed result in misaligned image-irradiance pairs.. Since the date modified seems much better aligned with the GHI time series, we use it to label the images in $\mathcal{D}^{\text{folsom}}$, except during the experiment in subsubsection 5.2.1, where we provide a more quantitative analysis on this issue.
@@
As a precautionary measure, we also compared the metadata of $x^{\text{sirta}}$ and $x^{\text{nrel}}$ to their file names, and found that the NREL images also contain an irregularity. However, the date modified issue in this dataset seems much more complex than in $\mathcal{D}^{\text{folsom}}$. In some cases, aligning the images with the date modified yielded a more sensible irradiance label, while in others, the opposite was true. This led us to choose the most simple solution possible, in which we discard all samples where $|\text{Timestamp}(\text{Date Modified})-\text{Timestamp}(\text{File Name})|>30\text{ s}$. This processing step retains 65,202 of the original 66,908 image/irradiance pairs and ensures a more proper alignment for all retained pairs.
Another thing to keep in mind when aligning the images with the irradiance data is the irregularity in the image timestamps, which is especially true assuming the date modified as the correct timestamp (see Figure 8). Since the irradiance samples are all synchronized at :00 seconds, aligning an image that was captured at :30 seconds with the nearest irradiance sample could result in a large error, especially in fast-moving cloud conditions. To mitigate this problem, we linearly interpolate the irradiance samples before the alignment. While this mostly affects $\mathcal{D}^{\text{folsom}}$, the linear interpolation is applied to the other two datasets as well.
Figure 8: Image timestamp second counts in $\mathcal{D}^{\text{folsom}}$. The upper axis shows that the file name timestamps are mostly synchronized at either :$0$ or :59 seconds (i.e. the file ends in “$0$.jpg” or “59.jpg”). On the other hand, the lower axis shows that the date modified timestamps are uniformly spaced throughout the 60 possible values.
@@ 3.7 Sun Mask Generation
In subsubsection 5.2.6 we conduct an ablation experiment where we use a Sun mask as an additional input to the estimation model, which has shown to significantly benefit the performance in [22]. First, we compute the distance of the Sun center from the center of the image plane, denoted by R, which is a function of the focal length of the camera f and the angle from the optical axis $\Phi$. This mapping function will vary, since different types of lenses will have different projection types [52]. For the stereographic projection, the mapping function is $\text{R}=2f\tan{\frac{\Phi}{2}}$, and for the equidistant projection the mapping function is $\text{R}=f\Phi$, where $\Phi$ is replaced by the zenith angle $\theta_{z}$. Following this, we compute the Sun center coordinates on the image plane ($\text{x}_{c}$, $\text{y}_{c}$) according to:
@@ 3.8 Train, Test and Validation Splits
The remaining years of each data set are used for the train set. In order to optimize the hyperparameters of the model, we split the full train set into a smaller train set and a validation set using $K$-Fold cross validation (CV). More specifically, we use stratified group $K$-Fold (with $K=5$), where each day is assigned to a different group (to avoid leaking samples from the same day in the training/validation sets) and each bin in the original train set (see Figure 4) is assigned to a different class to maintain the original irradiance distribution in all 5 folds. After optimizing the hyperparameters of each model, we use the full train set to train a final model, which will be evaluated on the test set. Table 3 shows the number of samples and year of each set, after the application of all the pre-processing steps.
Table 3: Number of samples and year of each set after the application of all the pre-processing steps.
Train Test
Dataset All Year All Clear Cloudy Year
Folsom 426,515 2014-2015 220,755 117,423 103,332 2016
SIRTA 297,774 2017-2018 103,926 18,447 85,479 2019
NREL 48,329 2020-2022 16,873 6,973 9,900 2023
@@ 4 Methods
Our study investigates a two-step forecasting approach, akin to SkyGPT [20], depicted in Figure 10. A compelling aspect of this approach is that both modules can be designed and optimized entirely independently, i.e., the approach completely decouples future event prediction and GHI estimation. Furthermore, this approach allows us to obtain a lower bound on the GHI forecasting error, simply by replacing the predicted future images with the ground truth future images. Our experiments target primarily the GHI estimation module, however we also include proof of concept experiments where we compare the two-step and single-step forecasting approaches depicted in Figure 10.
In subsection 4.1, we detail our training setup, the model architectures we trained and the evaluation metrics used in the GHI estimation experiments. All of the GHI estimation models are trained and evaluated separately for each dataset, however we also evaluate the aggregated performance across all three datasets.
In subsection 4.2, we briefly describe our training objectives and implementations details for the forecasting experiments, which are only evaluated on $\mathcal{D}^{\text{sirta}}$.
Figure 10: Two-step and single-step approaches for GHI nowcasting. Given an ASI sequence input of $T$ frames, each frame is transformed into a sequence of $N$ non-overlapping patches. In the two-step approach, these patch sequences will be fed into a video prediction module that predicts future ASI frames, and a subsequent module estimates the corresponding irradiance value for each generated frame. It is important to note that this irradiance estimation module is trained independently on the ground truth images. In the single-step approach, the irradiance forecasting module directly predicts the future irradiance values, by applying a linear regression head to the output of the learnable patch embedding, which we prepend to all $T$ patch sequences.
4.1 GHI Estimation
4.1.1 Training Setup
Given an input image $x_{i}$, each model produces an estimate $\hat{I}_{i}$ of the concurrent GHI measurement $I_{i}$. The models are trained to minimize the Mean Squared Error (MSE) loss function given by:
@@ 4.1.1 Training Setup
Our models are trained with a learning rate that exponentially decays from its initial value $\text{lr}_{0}$ to $\frac{\text{lr}_{0}}{10}$ during the first 75% of training according to:
@@
where n_epochs is the total number of epochs and $0<i<0.75\times\text{n\_epochs}$. The remaining 25% of training is done with the learning rate fixed at $\frac{\text{lr}_{0}}{10}$ and we start averaging the weights of the model at the end of every epoch, as this has shown to lead to better generalization [53]. Each model is trained on a NVIDIA GeForce RTX 3090 GPU (24 GB memory) for a total of 16 epochs, except during hyperparemeter optimization, where they are only trained for 8 epochs to save on computational time. Details on the initial learning rate and other hyperparameters values are described in the following section.
4.1.2 Model Architectures and Hyperparameter Optimization
@@
We search 30 different hyperparameter combinations per model per dataset, totaling 900 different searches, which are made through Bayesian optimization [57] using the train and validation splits defined in subsection 3.8 and with the goal of minimizing the average validation loss across all 5 folds. For hyperparameter optimization, we reduce the number of training epochs to 8 and only use 10% of the training data in each dataset. The hyperparameter search space is shown in Table 4.
@@
Initial Learning Rate ($\text{lr}_{0}$) $1,2,3,4,5,6,7,8,9,10]\times 10^{-4}$
@@
4.1.3 Evaluation Metrics
A large variety of metrics have been proposed by the solar forecasting community in order to assess the quality of a model’s predictions in the most insightful way possible [60, 61, 62]. In this study, we evaluate the root mean squared error (RMSE), the mean absolute error (MAE) and the normalized RMSE (nRMSE), which are well established metrics for evaluating the overall fit of the predictions:
@@ 4.1.3 Evaluation Metrics
$\text{nRMSE}=\frac{\text{RMSE}}{\frac{1}{N}\sum_{i=1}^{N}I_{i}}$ (16)
where $N$ is the number of samples, $I_{i}$ is the ground truth measurement of the GHI for the $i^{th}$ sample and $\hat{I}_{i}$ is the model’s estimate of the GHI for the $i^{th}$ sample. We employ the RMSE as our main ranking metric, since it penalizes larger errors more heavily than the MAE. The nRMSE is used specifically to analyze performance across different weather seasons and hours of the day, due to the scale-invariant nature of this metric.
Note that, regardless of our choice for the model’s target variable, the evaluation metrics are always expressed in terms of the GHI ($\text{W/m}^{2}$) or as a percentage of GHI (in the case of nRMSE), as that is the variable of interest.
4.2 Forecasting
In the forecasting task, we use $\mathcal{D}^{\text{sirta}}$ with the goal of generating future GHI values with lead times ranging from 2 to 10 minutes, based on 8 minutes of past ASI context. We investigate two approaches: a single-step approach, where an irradiance forecasting module directly predicts the future irradiance values, and a two-step approach, where a video prediction module first generates future ASI frames, and a subsequent module estimates the corresponding irradiance values for each generated frame. This irradiance estimation module corresponds to our best GHI estimation module, which is trained on the ground truth images.
4.2.1 Data Processing
In order to create the ASI sequence samples, we closely followed the data processing pipeline in [40], which is used as the baseline for our forecasting results. This lead to a few modifications to the data processing steps used in our GHI estimation experiments, described in section 3. Briefly, the key differences are: 1) the test set only includes samples that correspond to the odd days of 2019; 2) each image in the input sequence has a $128\times 128$ resolution.
Regarding the use of only odd days of 2019 as the test set, this only means that the test set in the GHI forecasting experiments will be about half the size of the test set in the GHI estimation experiments. We note, however, that the training set years remain the same, which means that, for the two-step approach, there is no data leakage between the video prediction and the GHI estimation modules.
Regarding the use of $128\times 128$ image resolution, this means that we need to resize the generated images to $64\times 64$ prior to the GHI estimation module in the two-step approach.
4.2.2 Two-step Approach
Given an input batch of image sequences $X_{i}\in\mathbb{R}^{B\times T\times C\times H\times W}$, the model is trained to predict the output image sequence $Y_{i}\in\mathbb{R}^{B\times T\times C\times H\times W}$ by minimizing the pixel-wise MSE, where $B$ is the batch size, $T$ is the number of frames, $C$ is the number of image channels, $H$ is the image height and $W$ is the image width. We use $T=5$ frames, such that the input consists of frames at times $\{t-8,t-6,t-4,t-2,t\}$, and the model predicts frames at $\{t+2,t+4,t+6,t+8,t+10\}$, in minutes. Our batch size is set to 16.
We employ an encoder-only Vision Transformer, closely following the design of PredFormer [63], which transforms the input image sequence $X_{i}$ into a sequence of $T\cdot N$ non-overlapping patches, where $N=\frac{HW}{p^{2}}$ is the number of patches per frame and $p$ is the patch size. We use a patch size of 16, resulting in $N=64$ patches per frame. Each image patch gets transformed into a patch embedding of dimension $D=256$, resulting in a $X_{p}\in\mathbb{R}^{B\times T\times N\times D}$ tensor.
A positional encoding is then added to $X_{p}$ to provide both the spatial and temporal information of each patch. Following this, each patch embedding will go through the stack of PredFormer layers, before finally being decoded back into the image patches by the linear patch decoder layer, which decodes $X_{p}$ into $Y_{i}$.
Regarding the PredFormer layers, we use full spatial-temporal self-attention, which allows each query patch to attend to all other patch tokens across the input sequence. Under this setup, each query patch will predict the patch at the exact same spatial position, but 10 minutes into the future, as illustrated by Figure 11. For more details on the model architecture, we refer the reader to the PredFormer article [63].
Finally, to obtain the GHI predictions, each of the $T$ predicted frames in $\hat{Y}_{i}$ are resized to a $64\times 64$ resolution, and are then fed to our best pre-trained estimation model to obtain the corresponding GHI estimate.
Figure 11: Input patches (left) and output patches (right) of the vision transformer model. Each query patch (red) can attend to all input patch tokens through full spatial temporal self-attention (blue). Each patch has size $16\times 16$, however, for better visualization, we show patches of size $32\times 32$ above.
4.2.3 Single-step Approach
This approach is conceptually similar to the first step of the two-step approach, and also follows the design of Predformer. However, instead of training the model to predict the future image frames, it is trained to directly predict the output clear sky index sequence $k_{t,i}\in\mathbb{R}^{B\times T}$ by minimizing the MSE
$L_{\text{MSE}}=\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{t=1}^{T}\left(k_{t,i}-\hat{k}_{t,i}\right)^{2}$ (17)
where $B$ and $T$ have the exact same values used in the first step of the two-step approach. To optimize the PredFormer model with this loss function, two other changes were made. First, we prepend a learnable token embedding $e\in\mathbb{R}^{1\times 1\times 1\times D}$ to every frame of every batch sample in $X_{p}$, such that the new $X_{p}\in\mathbb{R}^{B\times T\times(N+1)\times D}$. This is analogous to the CLS token used in BERT [64]. Second, the linear patch decoder layer is replaced with a linear irradiance regression head. Again, akin to BERT, this linear regression layer is only applied to the output of the $e$ token, obtaining the clear sky index prediction, which we transform into a GHI estimate through Equation 2. Everything else is the same as in the first step of the two-step approach.
4.2.4 Evaluation Metrics
We evaluate the results at every timestep $\{t+2,t+4,t+6,t+8,t+10\}$ using both the RMSE and the forecasting skill (FS). The forecasting skill is a metric that measures the overall improvement over the SPM, according to:
$\text{FS}=\left(1-\frac{\text{RMSE}}{\text{RMSE}_{\text{SPM}}}\right)$ (18)
where $\text{RMSE}_{\text{SMP}}$ is the error of the SPM predictions, given by Equation 3. A positive FS implies that a model is better than the SPM and a negative value suggests otherwise.
@@ 5 Experiments and Results
5.1 Benchmarking Results
Table 5 shows the RMSE and MAE for the test set for all ten models and all three datasets, as well as the aggregated performance across all datasets. From these results, we observe that model depth and the usage of residual connections improves performance. For example, the ResNet50 and EfficientNet models, which are both deep residual models, consistently outperformed the others, especially under cloudy sky conditions. This is expected, since deeper networks are better equipped to extract complex spatial features, which are crucial under these sky conditions, due to the high variety of shapes, sizes and possible overirradiance effects that clouds can have.
However, models that are deep but lack residual connections, such as SolarNet and VGG16, showed significantly worse performance. This aligns with the findings of [10], who demonstrated that residual connections facilitate the information and gradient flow through the network, allowing for a much more efficient training of very deep models. The shallowest models in our benchmark, Sunset and UNet, yielded the worst performance. Their limited depth is likely a determinant factor in this poor performance, since it restricts their ability to extract the complex spatial features of clouds.
Finally, the results show that $\mathcal{D}^{\text{folsom}}$ achieved the best “All” performance, despite yielding worst results on both the “Clear” and “Cloudy” subsets (an instance of Simpson’s paradox). The best “All” performance can be attributed to the fact that over 53% of the samples in $\mathcal{D}^{\text{folsom}}$ are clear sky samples (as depicted in Table 3), which will naturally bias the overall error toward lower values. On the other hand, the performance on the other two subsets was much worse, even compared to $\mathcal{D}^{\text{nrel}}$, which has far fewer training samples. We hypothesize that this behavior is due to the lens contamination we highlighted in subsection 3.1 and Figure 3, since the model can be mistaking the dirt stains for false clouds, hurting the performance on both the sky conditions.
@@ 5.1 Benchmarking Results
NREL EfficientNet 42.44 ± 4.58 52.14 ± 1.64 19.21 ± 13.28 23.97 ± 1.25 33.43 ± 1.11 10.53 ± 1.57
MobileNetV2 47.18 ± 1.18 60.57 ± 1.49 13.31 ± 0.53 27.15 ± 0.58 39.36 ± 0.77 9.80 ± 0.33
RegNet 41.65 ± 0.33 53.01 ± 0.44 14.44 ± 0.49 24.42 ± 0.17 33.47 ± 0.19 11.58 ± 0.29
ResNet18 40.84 ± 0.57 52.36 ± 0.75 11.91 ± 1.76 23.17 ± 0.54 33.12 ± 0.37 9.03 ± 1.25
ResNet34 43.12 ± 3.10 53.56 ± 3.19 20.25 ± 4.87 26.05 ± 2.57 34.00 ± 2.42 14.77 ± 3.11
ResNet50 38.75 ± 0.29 49.85 ± 0.37 10.25 ± 0.58 21.73 ± 0.23 31.56 ± 0.38 7.77 ± 0.41
SolarNet 44.33 ± 0.83 56.84 ± 0.98 12.93 ± 1.07 25.58 ± 0.84 36.62 ± 0.79 9.92 ± 1.08
Sunset 55.47 ± 2.76 71.24 ± 3.46 15.46 ± 1.42 31.14 ± 1.30 45.40 ± 1.79 10.89 ± 0.66
UNet 52.21 ± 1.77 66.45 ± 2.30 18.05 ± 1.07 31.43 ± 0.96 44.35 ± 1.60 13.09 ± 0.87
VGG16 41.45 ± 0.95 53.19 ± 1.19 11.80 ± 0.81 23.62 ± 0.69 33.76 ± 0.73 9.24 ± 0.75
Aggregated EfficientNet 38.53 ± 0.31 49.07 ± 0.12 14.13 ± 1.48 21.89 ± 0.22 30.31 ± 0.19 10.17 ± 0.30
MobileNetV2 43.95 ± 1.34 55.19 ± 1.24 19.39 ± 2.71 26.57 ± 1.25 35.13 ± 0.99 14.66 ± 1.86
RegNet 40.33 ± 0.86 51.45 ± 0.90 14.37 ± 1.39 23.24 ± 0.73 32.29 ± 0.70 10.64 ± 0.83
ResNet18 41.74 ± 0.67 52.98 ± 0.88 16.17 ± 0.26 24.45 ± 0.46 33.34 ± 0.68 12.08 ± 0.22
ResNet34 45.41 ± 8.03 51.12 ± 0.46 31.40 ± 21.50 24.17 ± 1.30 31.50 ± 0.29 13.98 ± 3.00
ResNet50 38.05 ± 0.20 48.75 ± 0.24 12.51 ± 0.36 21.39 ± 0.14 29.95 ± 0.17 9.48 ± 0.22
SolarNet 43.23 ± 1.36 54.56 ± 1.21 17.96 ± 2.59 25.30 ± 1.06 33.97 ± 0.79 13.25 ± 1.45
Sunset 48.14 ± 1.35 59.94 ± 1.44 23.15 ± 3.06 29.35 ± 1.23 38.16 ± 0.86 17.10 ± 2.27
UNet 49.54 ± 1.35 59.18 ± 0.97 31.29 ± 4.41 32.58 ± 1.78 38.96 ± 0.79 23.71 ± 3.60
VGG16 44.08 ± 1.06 55.61 ± 1.11 18.50 ± 1.42 26.22 ± 0.78 35.64 ± 0.79 13.11 ± 0.79